# Rethinking the Relationship between the Power Law and Hierarchical Structures

### v2

@@ 3.4 Context-free Independence Breaking
The CFIB is defined for the distribution of trees, as is the MI. We can estimate this metric from the dataset of trees, similarly to the estimation of the MI. We sample $N_{\text{data}}$ pairs of nodes that are tagged as $x_{0}$ and $x_{1}$, with the structural distance of $r_{\text{str}}$. Then, we compute the empirical distributions of tags on the children of the nodes. From these distributions, we estimate $J$ in the same procedure explained in the previous section. We also distinguish between pairs and their permutations, as in the estimation of MI.
We verify the accuracy of the estimates for the CFIB by varying $N_{\text{data}}$, as we did for the MI. Appendix C also presents the results obtained with alternative estimators, further supporting the validity of our estimates.
@@ 3.5 PCFG Approximation
We also construct a PCFG that approximates the distribution of natural language parse trees. This allows for a more detailed comparison between syntactic structures and PCFGs. The PCFG is constructed using the maximum likelihood estimation, which is equivalent to simply counting the frequency of each production rule in the dataset. Using the PCFG derived from the preprocessed dataset from BLLIP, we generate trees $1.6\times 10^{7}$ times, with up to $1\times 10^{2}$ iterations per generation.This procedure yields approximately 13 million terminated trees. We then apply the same analyses to these generated trees as to parse trees.
@@ 3.6 Fitting
We test whether the decay of correlation follows an exponential or power-law form by fitting the data with an exponential function, $A\exp(-\lambda r)$, and a power-law function, $Br^{-\alpha}$. To perform the fitting, we apply the least squares method after log-scaling the MI and CFIB values, because fitting with a linear scale puts too much weight on data points at shorter distances. We also fit the growth of sequential distance using an exponential function, $C\exp(\mu r)$, and a power-law function, $Dr^{\beta}$, without log-scaling the values. For both fitting, we use the lmfit Python package Newville et al. (2014). In the following results, we present the fitted parameters $A$, $\lambda$, and others along with the reduced chi-square statistic, $\chi^{2}_{\nu}$.
@@ 4.2 Mutual Information in Structures
Model fitting reveals that the correlation decays according to a power law ($\chi^{2}_{\nu}=1.1\times 10^{-1}$) rather than an exponential form ($\chi^{2}_{\nu}=1.3$) with respect to the structural distance. This result indicates that syntactic structures do not meet the condition necessary for the previously proposed argument to hold.
@@ 4.3 Growth of Sequential Distance
Question (ii) asks whether the sequential distance grows exponentially with increasing structural distance. To examine this, we count the number of POS tag pairs such that the structural and sequential distances are $r_{\text{str}}$ and $r_{\text{seq}}$, respectively. The results are shown in Figure 7. We also compute the average sequential distance for each structural distance, represented by black circles, and fit the results using both exponential and power-law functions.
The fitting reveals that the growth follows a power law ($\chi^{2}_{\nu}=1.4\times 10$) rather than an exponential form ($\chi^{2}_{\nu}=1.4\times 10^{21}$), although the trend becomes less clear for larger structural distances $r_{\text{str}}\gtrsim 50$ because of data sparsity. Notably, the power law exponent $\beta=0.72$ is smaller than $1$, indicating that the growth is slower than linear. As demonstrated in Figure 1 (b), this slow growth quantitatively reflects a strong branching bias in syntactic structures. These findings clearly show that the second condition does not hold for syntactic structures.
@@ 4.4 Context-free Independence Breaking
Question (iii) concerns whether the statistical properties of syntactic structures deviate significantly from those of PCFGs. We measure the CFIB $J$ as a metric of this deviation. In Figure 8, we present the results for fixed nodes being $(x_{0},x_{1})=(\text{NP},\text{NP})$, the most frequent pair in the corpus. The convergence of the estimates with increasing $N_{\text{data}}$ confirms that the estimation bias is negligible.
The CFIB values are clearly positive, meaning that syntactic structures do not have the essential property of PCFGs. Furthermore, the metric decays according to a power law ($\chi^{2}_{\nu}=8.1\times 10^{-2}$) rather than an exponential form ($\chi^{2}_{\nu}=7.1\times 10^{-1}$) with respect to structural distance, suggesting that the context-free independence is significantly broken even at large distances. These results indicate that the answer to question (iii) is negative.
@@
As shown in Appendix D, similar behaviors are observed for other frequent node pairs $(x_{0},x_{1})=(\text{NP},\text{VP})$ and $(x_{0},x_{1})=(\text{VP},\text{VP})$, suggesting that the observed deviation from PCFGs is not specific to a particular syntactic category pair.
@@ 4.5 Other Corpora
Figure 9 summarizes the results. For both corpora, the statistical properties are similar to those observed for BLLIP: the MI decays according to a power law in both sequences and structures; the growth of sequential distance is slower than linear; and the CFIB also decays following a power law. These behaviors are supported by the chi-square statistics, which are smaller for power-law fitting than for exponential fitting across all statistical quantities.
The results for these corpora also exhibit slight differences from those for BLLIP. $I_{\text{POS}}$ and $I_{\text{tag}}$ for WikiText, as well as $I_{\text{POS}}$ for NPCMJ, increase at large distances. This effect may be attributable to structural differences in unusually long sentences or to potential parsing errors in such cases. However, we cannot determine the significance of this behavior because the datasets contain too few long sentences.
Nonetheless, the results clearly support that the assumptions proposed by Lin and Tegmark (2017) are violated not only for BLLIP, but also for another English corpus, WikiText, as well as a Japanese treebank, NPCMJ. These results imply the robustness of our conclusion, given that Japanese is typologically distinct from English, exhibiting head-final, left-branching syntactic structures.
@@ 4.6 PCFG Approximation
The MI in structures, shown in Figure 10 (b), exhibits a power-law decay similar to that in syntactic structures, although the decay rate differs. This result does not contradict the properties of PCFGs proven by Lin and Tegmark (2017). While the MI in structures generated by PCFGs asymptotically follows an exponential decay at sufficiently large distances, the MI can exhibit different behavior at shorter distances. The non-exponential behavior of $I_{\text{tag}}$ suggests that for individual sentences, the structural distances are too short for the asymptotic behavior to emerge.
If the generation process continued for more steps, producing sufficiently large trees, the MI in Figure 10 (b) would converge to an exponential decay at large structural distances. However, such a setup cannot be justified because the generated trees would be much larger than typical parse trees. Even if we assume this setup, the MI in sequences would decay more rapidly than a power law as long as the trees are sufficiently biased so that the sequential distances grow more slowly than exponentially. The mechanism of the power-law decay of correlation observed in natural languages remains unexplained.
The non-exponential behavior of $I_{\text{tag}}$ in the PCFG has another important implication. As discussed in Section 2.2, the argument by Lin and Tegmark (2017) can, in principle, be extended to hierarchical structures beyond syntax. Following this perspective, several studies have applied the argument to domains other than texts and speeches by human adults. These studies reported the power-law decay of correlation in child speeches and animal signals, interpreting it as evidence for the existence of underlying hierarchical structures in these domains (Sainburg et al., 2019; Sainburg and Gentner, 2021; Sainburg et al., 2022; Howard-Spink et al., 2024; Youngblood, 2024).
However, the argument by Lin and Tegmark (2017) relies on the asymptotic behavior of PCFGs, which does not appear when the observed sequences are too short. In child speeches and animal signals, the sequences are typically shorter than sentences produced by human adults. In such cases, it is difficult to interpret the power-law decay of correlation as evidence for hierarchical structures.
@@
In Figure 10 (d), the estimates of CFIB is much smaller than those for syntactic structures in Figure 8. Furthermore, the estimates continue to decrease in the log-scale plot as $N_{\text{data}}$ increases, which is the expected behavior. This result is in contrast to that for syntactic structures in natural languages.
@@ 5 Discussion and Conclusion
The present study examines this relationship, specifically the argument proposed by Lin and Tegmark (2017). They attributed the power-law decay of correlation in natural languages to underlying hierarchical structures, which has since inspired numerous studies on both human languages and animal signals. However, their argument relies on several assumptions that are questionable from a linguistic perspective. To examine the validity of the assumptions for syntactic structures, we test the three key questions through statistical analyses of parse trees in natural languages.
The analyses yield negative answers to all three questions: (i) The correlation in syntactic structures decays according to a power law with respect to the structural distance; (ii) The growth of the sequential distance is slower than linear with the structural distance, due to the strong branching bias in syntactic structures; (iii) The statistical properties of syntactic structures differ from those of PCFGs, particularly in the CFIB. These results reveal that the relationship between the power-law decay of correlation and syntactic structures differs from that proposed by Lin and Tegmark (2017).
We obtained similar results under various settings: for the English treebank BLLIP with several different schemes, for another English corpus, WikiText, and for the Japanese treebank NPCMJ. The consistency of the results across different schemes, domains, and languages suggests the robustness of our conclusion.
We also analyze the PCFG obtained from the maximum likelihood estimation of BLLIP. The statistical properties of the PCFG clearly differ from those in parse trees. Furthermore, this analysis demonstrates that the exponential decay of $I_{\text{tag}}$ in the PCFG does not emerge within individual sentences. Although Lin and Tegmark (2017) proved that $I_{\text{tag}}$ decays exponentially at sufficiently large structural distances, the structural distances in typical parse trees are too short for this behavior to appear. This result highlights the difficulty of applying the proposed argument not only to sentences produced by human adults but also to child speeches and animal signals, in which the sequences are shorter than sentences by human adults.
@@
Currently, we cannot rule out the possibility that the MI in natural language syntactic structures follows an exponential decay at structural distances much larger than those observed in this study. However, even if this is the case, such large parse trees would be too rare to have a significant influence on the statistical properties of the corpus as a whole. Moreover, as long as syntactic structures are biased, the decay of MI in sequences should be faster than a power law, because the sequential distance grows more slowly than exponentially. Therefore, statistical analyses of larger parse trees are unlikely to reveal the mechanism underlying the power-law decay of correlation in natural languages.
@@ Acknowledgments
We thank Taiga Ishii, Shinnosuke Isono, Sho Yokoi, and Keisuke Sakaguchi for useful discussions and suggestions. This work was supported by JSPS KAKENHI Grant Nos. 23H01095, 23KJ0622, 24H00087, and 25K24434; JST, PRESTO Grant No. JPMJPR21C2; and the World-Leading Innovative Graduate Study Program for Advanced Basic Science Course at the University of Tokyo.
Appendix A Tag Reduction
To reduce the number of categories, we grouped tags with similar functions into the same category. The eleven categories used for BLLIP and WikiText and the seven categories for NPCMJ are shown in Tables 1 and 2, respectively.
@@ Appendix A Tag Reduction
Appendix B Other Schemes
@@
Although the statistical quantities depend on the specific preprocessing scheme, we expect that these details do not affect our conclusions, such as whether the MI decays according to a power law or an exponential function, and whether the CFIB is significantly positive. To examine this expectation, we perform additional analyses under two alternative preprocessing schemes: the unbinarized setting and the phrasal-tag-only setting, in which POS and phrasal tags are distinguished and only the latter are analyzed.
Figure 11: (a1, 2) Results for the unbinarized setting: (a1) MI $I_{\text{tag}}$ between POS or phrasal tags, where fitting was performed for $N_{\text{data}}=2.56\times 10^{6}$; (a2) distribution and average of sequential distances. (b1, 2) Results for the phrasal-tag-only setting: (b1) MI $I_{\text{tag}}$ between phrasal tags, where fitting was performed for $N_{\text{data}}=2.56\times 10^{6}$; (a2) CFIB $J$, where fitting was pwerformed for $N_{\text{data}}=6.4\times 10^{5}$.
Figures 11 (a1) and (a2) show the results for the unbinarized setting. Because binarization does not affect the MI in sequences and the dataset is too limited to reliably estimate the CFIB for unbinarized trees, we present only the MI in structures and the sequential distance as a function of the structural distance. For both quantities, the results are qualitatively similar to those for the binarized setting in the main text: the decay of $I_{\text{tag}}$ follows a power-law function rather than an exponential function, and the growth of $r_{\text{seq}}$ is slower than exponential.
@@
Across both alternative preprocessing schemes, our main conclusion holds: the assumptions in the argument by Lin and Tegmark (2017) are not satisfied. These results indicate that our findings are robust to the details of preprocessing.
Appendix C Alternative Estimators

### v3

@@ 3.4 Context-free Independence Breaking
The CFIB is defined for the distribution of trees, as is the MI. We can estimate this metric from the dataset of trees, similarly to the estimation of the MI. We sample $N_{\text{data}}$ pairs of nodes that are tagged as $x_{0}$ and $x_{1}$, with the structural distance of $r_{\text{str}}$. Then, we compute the empirical distributions of tags on the children of the nodes. From these distributions, we estimate $J$ in the same procedure explained in Section 3.2. We also distinguish between pairs and their permutations, as in the estimation of MI.
We verify the accuracy of the estimates for the CFIB by varying $N_{\text{data}}$, as we did for the MI. Appendix D also presents the results obtained with alternative entropy estimators, further supporting the validity of our estimates.
@@ 3.5 PCFG Approximation
We also construct a PCFG that approximates the distribution of natural language parse trees. This allows for a more detailed comparison between syntactic structures and PCFGs. The PCFG is constructed using the maximum likelihood estimation, which is equivalent to simply counting the frequency of each production rule in the dataset. Using the PCFG derived from the preprocessed dataset from BLLIP, we generate trees $1.6\times 10^{7}$ times, with up to $100$ iterations per generation. This procedure yields approximately 13 million terminated trees. We then apply the same analyses to these generated trees as to parse trees.
@@ 3.6 Fitting
We characterize the decay of correlation by fitting the data with exponential and power-law functions. Our goal is not to fit the data as accurately as possible or to estimate parameter values, but to test whether the observed behavior is closer to a power law or to an exponential form. For this purpose, using simpler models with fewer parameters is more appropriate, as it avoids issues such as overfitting. Accordingly, we use the two simple models, $A\exp(-\lambda r)$ and $Br^{-\alpha}$.
To perform the fitting, we apply the least squares method after log-scaling the MI and CFIB values, because fitting with a linear scale puts too much weight on data points at shorter distances.
In model selection, model complexity should be taken into account, because goodness of fit, such as likelihood or squared error, trivially improves as the number of parameters increases. However, in our case, both models have two parameters. Therefore, it is sufficient to compare goodness of fit. Specifically, we use the reduced chi-square statistic, $\chi^{2}_{\nu}$. Under the commonly used assumption of Gaussian noise, comparing this statistic is equivalent to using standard model selection criteria such as the Akaike information criterion Akaike (1974) and the Bayesian information criterion Schwarz (1978).
We also fit the growth of sequential distance using an exponential function, $C\exp(\mu r)$, and a power-law function, $Dr^{\beta}$, without log-scaling the values.
For fitting, we use the lmfit Python package Newville et al. (2014). In the following results, we present the fitted parameters $A$, $\lambda$, and others along with the reduced chi-square statistic, $\chi^{2}_{\nu}$.
@@ 4.2 Mutual Information in Structures
Model fitting reveals that the correlation decays according to a power law ($\chi^{2}_{\nu}=1.1\times 10^{-1}$) rather than an exponential form ($\chi^{2}_{\nu}=1.3$) with respect to the structural distance. This result indicates that syntactic structures do not meet the condition necessary for the proposed argument to hold.
@@ 4.3 Growth of Sequential Distance
Question (ii) asks whether the sequential distance grows exponentially with increasing structural distance. To examine this, we count the number of POS tag pairs such that the structural and sequential distances are $r_{\text{str}}$ and $r_{\text{seq}}$, respectively. The results are shown in Figure 7. We also compute the average sequential distance for each structural distance, represented by black circles, and fit the results using exponential and power-law functions.
The fitting reveals that the growth follows a power law ($\chi^{2}_{\nu}=1.4\times 10$) rather than an exponential form ($\chi^{2}_{\nu}=1.4\times 10^{21}$), although the trend becomes less clear for larger structural distances $r_{\text{str}}\gtrsim 50$ because of data sparsity. Notably, the exponent $\beta=0.72$ is smaller than $1$, indicating that the growth is slower than linear. As demonstrated in Figure 1 (b), this slow growth quantitatively reflects a strong branching bias in syntactic structures. These findings clearly show that the second condition does not hold for syntactic structures.
@@ 4.4 Context-free Independence Breaking
Question (iii) concerns whether the statistical properties of syntactic structures deviate significantly from those of PCFGs. We measure the CFIB $J$ as a metric of this deviation. Since reliable estimation of this metric requires a sufficiently large number of samples, we focus on the three most frequent cases, namely the fixed node pairs $(x_{0},x_{1})=(\text{NP},\text{NP})$, $(\text{NP},\text{VP})$, and $(\text{VP},\text{VP})$.
The result for $(\text{NP},\text{NP})$ is shown in Figure 8. The convergence of the estimates with increasing $N_{\text{data}}$ confirms that the estimation bias is negligible. The CFIB values are clearly positive, meaning that syntactic structures do not have the essential property of PCFGs. Furthermore, the metric decays according to a power law ($\chi^{2}_{\nu}=8.1\times 10^{-2}$) rather than an exponential form ($\chi^{2}_{\nu}=7.1\times 10^{-1}$) with respect to the structural distance, suggesting that the context-free independence is significantly broken even at large distances. These results indicate that the answer to question (iii) is negative.
@@
As shown in Appendix E, similar behaviors are observed for other frequent node pairs $(x_{0},x_{1})=(\text{NP},\text{VP})$ and $(x_{0},x_{1})=(\text{VP},\text{VP})$, suggesting that the observed deviation from PCFGs is not specific to a particular syntactic category pair.
@@ 4.5 Other Corpora
Figure 9 summarizes the results. For both corpora, the statistical properties are similar to those observed for BLLIP: the MI decays according to a power law in both sequences and structures; the growth of sequential distance is slower than linear; and the CFIB also decays following a power law. These behaviors are supported by the reduced chi-square statistics, which are smaller for power-law fitting than for exponential fitting across all statistical quantities.
The results also exhibit slight differences from those for BLLIP. $I_{\text{POS}}$ and $I_{\text{tag}}$ for WikiText, as well as $I_{\text{POS}}$ for NPCMJ, increase at large distances. This effect may be attributable to structural differences in unusually long sentences or to potential parsing errors in such cases. However, we cannot determine the significance of this behavior because the datasets contain too few long sentences.
Nonetheless, the results clearly support that the assumptions proposed by Lin and Tegmark (2017) are violated not only for BLLIP, but also for WikiText and for NPCMJ. These results imply the robustness of our conclusion, given that Japanese is typologically distinct from English, exhibiting head-final, left-branching syntactic structures.
@@ 4.6 PCFG Approximation
The MI in structures, shown in Figure 10 (b), exhibits a power-law decay similar to that in syntactic structures, although the decay rate differs. This result does not contradict the properties of PCFGs proven by Lin and Tegmark (2017). While the MI in structures generated by PCFGs asymptotically follows an exponential decay at sufficiently large distances, the MI can exhibit different behavior at shorter distances. The non-exponential behavior of $I_{\text{tag}}$ suggests that for individual trees, the structural distances are too short for the asymptotic behavior to emerge.
This pre-asymptotic behavior in the PCFG constructed from parse trees raises the possibility that the statistical behavior of parse trees themselves is also pre-asymptotic. If so, it also calls into question the applicability of the argument of Lin and Tegmark (2017) to natural language parse trees, since the argument relies on asymptotic behavior.
This observation also has an important implication about the applicability of their argument to domains beyond texts and speech by human adults. As discussed in Section 1, several studies have applied their argument to child speech and animal behavior, interpreting the power-law decay of correlations in these domains as evidence for the existence of some form of underlying hierarchical structures (Sainburg et al., 2019; Sainburg and Gentner, 2021; Sainburg et al., 2022; Howard-Spink et al., 2024; Youngblood, 2024).
However, in these domains, the sequence lengths are typically comparable to those in the PCFG studied here. In such cases, it is difficult to interpret the power-law decay of correlation as evidence for hierarchical structures, since the asymptotic behavior is unlikely to emerge.
@@
In Figure 10 (d), the estimates of CFIB is much smaller than those for syntactic structures in Figure 8. Furthermore, the estimates continue to decrease in the log-scale plot as $N_{\text{data}}$ increases, which is the expected behavior. These results are in contrast to those for parse trees.
@@ 5 Discussion and Conclusion
The present study examines this relationship, specifically the argument proposed by Lin and Tegmark (2017). They attributed the power-law decay of correlation in natural languages to underlying hierarchical structures, which has since inspired numerous studies on both human languages and animal behavior. However, their argument relies on several assumptions that are questionable from a linguistic perspective. To examine the validity of the assumptions for syntactic structures, we test the three key questions through statistical analyses of parse trees in natural languages.
The analyses yield negative answers to all three questions: (i) The correlation in syntactic structures decays according to a power law with respect to the structural distance; (ii) The growth of the sequential distance is slower than linear with the structural distance, due to the strong branching bias in syntactic structures; (iii) The statistical properties of syntactic structures significantly differ from those of PCFGs. These results reveal that the relationship between the power-law decay of correlation and syntactic structures differs from that proposed by Lin and Tegmark (2017).
We obtained similar results under various settings: for the English treebank BLLIP with several different schemes, for another English corpus, WikiText, and for the Japanese treebank NPCMJ. To further assess the robustness of our conclusions, it is important to examine statistical behavior under additional annotation schemes, domains, and languages. In particular, revealing the extent to which these behaviors are shared across languages is important from a linguistic perspective.
We also analyzed the PCFG obtained from the maximum likelihood estimation of BLLIP. The statistical properties of the PCFG clearly differ from those in parse trees. Furthermore, this analysis demonstrates that the asymptotic behavior of the PCFG does not emerge within individual sequences. This implies the difficulty of applying the proposed argument not only to sentences produced by human adults but also to child speech and animal behavior, where sequence lengths are comparable to those for the PCFG.
@@
It is noteworthy that the pre-asymptotic behavior in the PCFG constructed from parse trees does not serve for a definitive evidence against the validity of the argument by Lin and Tegmark (2017), because this behavior does not necessarily mean that parse trees themselves also exhibit pre-asymptotic behavior. Our main evidence instead come from the direct analyses of parse trees in Sections 4.14.5, which clearly show that the statistical properties of natural language parse trees for their typical sizes differ from those assumed in the proposed argument.
Currently, we cannot rule out the possibility that the MI in natural language syntactic structures follows an exponential decay at structural distances much larger than those observed in this study. However, even if this were the case, it would not explain the power-law decay of correlation in natural language sequences, because such large parse trees would be too rare to have a significant influence on the statistical properties of the corpus. Moreover, as long as syntactic structures are biased, the sequential distance grows more slowly than exponentially, resulting in a decay of MI in sequences that is faster than a power law. Therefore, statistical analyses of larger parse trees are unlikely to account for the power-law decay of correlation observed in natural languages.
@@ Acknowledgments
We thank the action editor, Carlos Gómez-Rodríguez, and the anonymous reviewers for their constructive and helpful feedback. We also thank Taiga Ishii, Shinnosuke Isono, Sho Yokoi, and Keisuke Sakaguchi for valuable discussions and suggestions. This work was supported by JSPS KAKENHI Grant Nos. 23H01095, 23KJ0622, 24H00087, and 25K24434; JST, PRESTO Grant No. JPMJPR21C2; and the World-Leading Innovative Graduate Study Program for Advanced Basic Science Course at the University of Tokyo.
Appendix A Tree Statistics
To provide intuition about both the effect of preprocessing on tree shapes and the differences between English and Japanese, we compute basic tree-shape statistics on the original parse trees in BLLIP and NPCMJ. Specifically, we consider the following statistics:
Depth.
The maximum path length from the root to any node in a tree, averaged over all parse trees.
Branching Factor.
The number of children of an internal node, averaged over all internal nodes. This value ranges from $1$ upward and equals $2$ for strictly binary-branching trees without unary nodes.
Proportion of Unary Nodes.
An internal node is unary if it has exactly one child. We report the proportion of unary nodes, defined as the number of unary internal nodes divided by the total number of internal nodes, excluding POS-to-terminal productions.
As shown in Table 1, the branching factor and the proportion of unary nodes are comparable between BLLIP and NPCMJ. In contrast, the depth for NPCMJ is smaller than that for BLLIP, suggesting that parse trees in NPCMJ tend to be flatter.
Corpus Depth Branching Prop. Unary
BLLIP $11.598(\pm 4.354)$ $1.527(\pm 0.904)$ $0.222$
NPCMJ $6.642(\pm 3.336)$ $1.547(\pm 1.282)$ $0.271$
Table 1: Tree-shape statistics computed on the original parse trees in BLLIP and NPCMJ. The depth and the branching factor are reported as mean $\pm$ standard deviation across sentences.
Appendix B Tag Reduction
To reduce the number of categories, we grouped tags with similar functions into the same category. The eleven categories used for BLLIP and WikiText and the seven categories for NPCMJ are shown in Tables 2 and 3, respectively.
@@ Appendix B Tag Reduction
Appendix C Other Schemes
@@
Although the statistical quantities depend on the specific preprocessing scheme, we expect that these details do not affect our conclusions, such as whether the MI decays according to a power law or an exponential function, and whether the CFIB is significantly positive.
To examine this expectation, we perform additional analyses on BLLIP under two alternative preprocessing schemes: the unbinarized setting and the phrasal-tag-only setting, in which POS and phrasal tags are distinguished and only the latter are analyzed.
In addition, we analyze unbinarized trees in NPCMJ. Because the original trees in NPCMJ are flatter than those in BLLIP, as shown in Appendix A, the effect of binarization could be larger.
Figure 11: (a1, 2) Results for BLLIP under the unbinarized setting: (a1) MI $I_{\text{tag}}$ between POS or phrasal tags, where fitting was performed for $N_{\text{data}}=2.56\times 10^{6}$; (a2) distribution and average of sequential distances. (b1, 2) Results for BLLIP under the phrasal-tag-only setting: (b1) MI $I_{\text{tag}}$ between phrasal tags, where fitting was performed for $N_{\text{data}}=2.56\times 10^{6}$; (a2) CFIB $J$, where fitting was pwerformed for $N_{\text{data}}=6.4\times 10^{5}$. (c1, 2) Results for NPCMJ under the unbinarized setting: (a1) MI $I_{\text{tag}}$ between POS or phrasal tags, where fitting was performed for $N_{\text{data}}=2.56\times 10^{6}$; (a2) distribution and average of sequential distances.
Figures 11 (a1) and (a2) show the results for BLLIP under the unbinarized setting. Because binarization does not affect the MI in sequences and the dataset is too limited to reliably estimate the CFIB for unbinarized trees, we present only the MI in structures and the sequential distance as a function of the structural distance. For both quantities, the results are qualitatively similar to those for the binarized setting in the main text: the decay of $I_{\text{tag}}$ follows a power-law function rather than an exponential function, and the growth of $r_{\text{seq}}$ is slower than exponential.
@@
The MI in structures and the sequential distance for unbinarized trees in NPCMJ are presented in Figures 11 (c1) and (c2). Again, the results are similar to those in other cases. Both quantities display power-law dependence on the structural distance rather than exponential behavior.
Across all alternative preprocessing schemes, our main conclusion holds: the assumptions in the argument by Lin and Tegmark (2017) are not satisfied. These results indicate that our findings are robust to the details of preprocessing.
Appendix D Alternative Estimators